OpenStudy (anonymous):

How to solve for v in this fraction? putting it down below. Help is appreciated (:

OpenStudy (anonymous):

\[\frac{ 5 }{ 6 } = \frac{ v-2 }{ 8 }\]

OpenStudy (anonymous):

we can say : 5*8=6v+6*(-2) K ?:)

OpenStudy (anonymous):



OpenStudy (anonymous):

would I now distribute 6 to (v-2)?

OpenStudy (anonymous):

Yep !

OpenStudy (anonymous):

where do I go after that? 5*8 = 6v - 12

OpenStudy (anonymous):

then we have : 40=6v-12 40+12=6v

OpenStudy (anonymous):

divide by 6?

OpenStudy (anonymous):

Yes !

OpenStudy (anonymous):

52 doesn't divide evenly into 6. Is it okay to have an answer of v = 8.6?

OpenStudy (anonymous):

Yep !!!

OpenStudy (anonymous):

okay! thanks so much for the help!

OpenStudy (anonymous):

You re welcome !:)